# Novel View Synthesis for Cinematic Anatomy on Mobile and Immersive Displays

### v1

@@ 3.3 Compressed Differentiable 3D Gaussian Splatting
With differentiable 3DGS, the object is described by a set of 3D Gaussians
@@
Each Gaussian is centered at $x\in\mathbb{R}^{3}$, and a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$ is used to describe its orientation and shape. A Gaussian has an opacity $\alpha\in[0,1]$, and a view-dependent color that is represented by a set of spherical harmonics (SH) coefficients. For each view, SH coefficients are multiplied with the view direction to obtain the color [25].
The 2D projection of a 3D Gaussian is a 2D Gaussian with covariance
$\Sigma^{\prime}=JW\Sigma W^{T}J^{T},$ (2)
where $W$ is the view transformation matrix and $J$ is the Jacobian of the affine approximation of the projective transformation. By computing the exponential decay of color and opacity across the 2D projection, at each pixel the color and opacity that is seen along the ray of sight through the 3D Gaussian can be reconstructed (see Fig. 2). A final pixel color $C$ is then computed by blending the contributions of all Gaussians in sorted order:
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (3)
Here, $N$ is the number of Gaussians affecting a pixel, and $c_{i}$, $\alpha_{i}$, respectively, are the view-dependent color of a Gaussian and its opacity, modulated by the exponential falloff from the projected Gaussian’s center point.
@@
While Zwicker et al. [61] model a 3D scalar field via a set of 3D Gaussians so that the field can be reconstructed sufficiently well, Kerbl et al. [25] optimize the position, shape, opacity and SH coefficients of each 3D Gaussian so that their rendering matches a set of initial images of the object. The optimization is performed via differentiable rendering, by taking into account the changes in pixel color due to changes of the 3D Gaussian parameters. In the optimization process, initially selected 3D Gaussians are removed (if no contribution), adaptively split and their shapes and appearance attributes are modified to minimize an image-based loss function.
Figure 2: Illustration of Gaussian splatting. Each 3D Gaussian has an opacity and a set of SH coefficients, to evaluate the per-pixel opacity and view-dependent color of the Gaussian’s 2D projection.
Figure 3: Test image reconstruction using different initialization schemes. Experiments were performed with Fullbody.
Figure 4: From left to right, Brain ($3224\times 3224\times 3585$), Kidney ($1510\times 1706\times 1415$) and Fullbody ($317\times 317\times 835$). All images rendered at full HD with HR-compressed 3D Gaussian splatting using $<$ 70 MB per dataset at $>$ 30 frames per second.
3.3.1 Compression
Due to the use of per-splat SH coefficients and shape attributes, the 3D Gaussian splat representation can become memory consuming and take up hundreds of MB or even GB per scene. Niedermayr et al. [38] have proposed strategies for compressing this representation. SH coefficients and Gaussian shape parameters are encoded into compact codebooks via sensitivity-aware vector quantization. To reduce the error introduced by vector quantization, the parameters are then fine-tuned on the training images after compression. To represent the scene parameters with fewer bits, quantization-aware training [41] is used. The final set of 3D Gaussians is linearized along a space-filling curve in the 3D domain. In this way, spatial coherence is maintained in the linearized sequence, and entropy encoding can be used to further compress the sequence of splat parameters.
For volume reconstruction, we have modified the compression scheme in the following way: Firstly, since Cinematic Anatomy requires the entire dataset in focus, we remove the scaling factor that is stored per Gaussian to represent scenes with objects in focus (small Gaussians) and surrounding background (large Gaussians). We call this strategy high-rate-compression (HR-compression). Since quantization of SH coefficients and shape parameters can introduce compression errors, we provide a variant that compress the final Gaussian representation at no perceivable loss in reconstruction quality. Therefore, we only use quantization-aware training to reduce all scene parameters but the Gaussians’ positions to an 8-bit representation during optimization. Positions are encoded using 16 bit floating point numbers. We will subsequently call this strategy high-quality-compression (HQ-compression). On average, it gives a compression factor of 26×26$\times$26 × compared to the uncompressed 3DGS representation, bringing all scenes bellow 170 MB.
3.3.2 Volume Guided Initialization
When using 3DGS, an initial set of 3D Gaussian kernels is first selected. These Gaussians are then removed, split or re-positioned, and the shape and appearance of the Gaussian kernels is optimized. Kerbl et al. [25] obtain the initial positions of the 3D Gaussians from the given images with structure from motion, or with random initialization were Gaussians are randomly positioned in the scene. For volume rendering, we randomly place Gaussians within the volume bounding box and set their initial color to grey. All other parameters are initialized as proposed by Kerbl et al. [25].
In Fig. 3 we exemplarily compare the effectiveness of the different initialization schemes based on optimization convergence for one of our test datasets. An initialization with the Gaussians’ positions and colors from a previous reconstruction is used as gold standard. As can be seen, while all initialization techniques reach the same level of fidelity, volume-guided initialization does so with less iteration steps. However, it is fair to say that in all of our experiments the performance improvements were overall not significant, so that we decided to use random initialization in all upcoming tests.
3.3.3 Alpha Channel Reconstruction
@@
In contrast to classical novel view synthesis, where only RGB colors are reconstructed, in volume rendering applications also the per-pixel accumulated opacity (i.e., alpha) needs to be reconstructed for blending correctly over the background. We extended 3DGS to allow for the differentiable rendering of images with alpha channel. Similar to Kerbl et al. [25] we use a combination per pixel L1 and SSIM Loss to faithfully reconstruct the alpha channel of the training images. This greatly improves the reconstruction quality as we show in Section 4.4.
3.3.4 Mip Splatting
Scenes rendered with 3DGS can show severe artifacts when novel camera perspectives diverge from those the 3D Gaussian representation was optimized for. Yu et al. [59] name the following two reasons for this behavior: Firstly, the 3D Gaussian representation exhibits frequencies that are too high to be faithfully reconstructed by the used sampling rate. Secondly, during splat-based rendering, a 2D dilation filter is applied that causes artefacts when zooming out and 2D splats become too small.
4 Results
@@
We analyze the performance, memory consumption and quality of the proposed pipeline for Cinematic Anatomy with a variety of high-resolution medical datasets showing different anatomical structures.
Our 3DGS implementation is a modification of the code implementation provided by Kerbl et al. [25]. For compression and rendering, we use the settings described in [38] and their publicly available WebGPU renderer.
@@ 4.1 Datasets
The hierarchical phase-contrast tomography (HiP-CT) data was acquired at the European Synchrotron Radiation Facility (ESRF) in the context of the Human Organ Atlas project[51] [^1] https://human-organ-atlas.esrf.eu.
@@
Fullbody is a human CT angiography scan at resolution $317\times 317\times 835$ from collection [52], image id s0287. The dataset contains some semi-transparent material that shows significant differences under directional lighting. It is rendered under complex lighting conditions to demonstrate reconstruction quality also in this situation.
All datasets are shown in Fig. 4. For each dataset, between one and three rendering presets have been used, which include segmentations, the selected transfer function and lighting conditions. 3DGS optimization has been performed training images of resolution $2048\times 2048$.
@@ 4.2 Preprocessing
With a GPU providing sufficient RAM, the initial images of all datasets can be generated with the publically available Cinematic Anatomy package [^2] https://siemens-healthineers.com/cinematic-anatomy, e.g. on an NVIDIA RTX 8000/A6000 GPU with 48GB RAM and using the built in animation system to generate the views. All images used as training and test data in our experiments have been rendering on a research version providing batch rendering support, running on an NVIDIA A100 GPU for the Brain data and NVIDIA RTX A5000 for the Kidney and Fullbody data. Note that rendering HipCT in its original resolution on a single GPU requires bricking the data into multiple smaller blocks. This increases the rendering times drastically, since path tracing requires streaming each block multiple times from the CPU to the GPU.
Table 1 shows in columns Size the size of each dataset in GB compared to the size of the final Gaussian representation in MB, when compressed using HR-compression. Column Views shows the number of training images used for differentiable Gaussian splatting optimization. Columns Time show the times to render the initial images via path tracing, and the computation times for generating the compressed Gaussian representations. Note that $90\%$ of the latter time are required by the optimization to generate the 3D Gaussian representation, and only about $10\%$ are consumed by the compression. Column Gaussians gives the number of 3D Gaussians in the final representation. Note that for the body scan, the used preset uses a lung mask-dependent and clip-plane dependent transfer function, which slows down the rendering considerably. As can be seen in columns Size, the compressed Gaussian representation is so small that it can be downloaded over low-bandwidth channels and rendered on mobile devices equipped with mid- or even low-end GPUs.
Table 1: Memory and preprocessing statistics using HR-compression. Timings given for $2048\times 2048$ training images.
Path Tracing 3DGS
@@
Fullbody 0.2 GB 23 Min 99 7 MB 50 Min 0.9 M
@@ 4.3 View Selection
Automatic view selection is demonstrated with Fullbody, which exhibits a lot of structures which are not visible from cameras placed on an ellipsoid around the volume. As a baseline we reconstruct the volume with images from 256 randomly placed cameras on the ellipsoid. For comparison, we reduce this number to 128, and generate 128 additional cameras with the proposed view selection algorithm (see Fig. 5). As can be seen, overall improved reconstruction quality of parts not seen with random camera selection is achieved.
Refer to captionRefer to caption
Figure 5: Our view selection algorithm generates cameras covering unseen parts of the volume like the inside of the rib cage. The renderings show the improvement in reconstruction.
@@ 4.4 Quality Evaluation
To support a qualitative analysis, Fig. 8 and Fig. 9 compare test images that have not been seen during 3DGS optimization to the same views rendered with HQ-compressed and HR-compressed 3DGS. Close-up views emphasize that only very subtle color shifts between path traced images and images generated via HR-compressed 3DGS can be observed. HQ-compression leads to an increase in memory of a factor of three, yet differences in image quality are further reduced and become so small that they are hardly noticeable by eye.
To quantify these differences, Table 2 shows the average SSIM and PSNR between the test images and the novel views rendered with HR-compressed 3DGS.
Table 2: Quantitative evaluation of HR-compression, using $2048\times 2048$ training images and averaged over all presets.
@@
Fullbody 0.87 26.90 29.57
For PSNR and SSIM only pixels which are not empty (alpha $>0$) in the rendered and ground truth image are considered. PSNR (Alpha) measures the PSNR for the alpha channel between rendered images and ground truth.
It is worth noting here, that significant losses in reconstruction quality are introduced when differentiable 3DGS is used for solely optimizing RGB color (see Fig. 6 for an example). As can be seen, extending 3DGS so that also opacity is considered in the optimization process improves greatly the reconstruction quality and removes unwanted artifacts caused by the background.
In a final experiment we shed light on the capabilities of 3DGS to reconstruct semi-transparent regions in a dataset. Fullbody, with a corresponding preset, is used here as an example. The preset has been selected so that certain tissue types in the dataset become semi-transparent. The bottom images in Fig. 9 show the application of this preset, including a test image and the corresponding novel views generated with compressed 3DGS. It can be seen that overall the novel view matches the test images fairly well. When looking at the closeup views, however, one sees that some small details are not reconstructed accurately, and that especially the semi-transparent structures are blurred out in the final images.
To further shed light on this situation, we have also experimented with a setting with strong directional lighting from the used environment map. In such a setting, one observes some high-frequency illumination variations especially in the volumetric regions, making it more difficult for 3DGS to accurately recover the structures. Interestingly, Fig. 7 demonstrates that the reconstruction of illumination variation works very well and does not show any severe reconstruction artefacts. At the same time, the semi-transparent regions are again blurred out to a certain extent. We believe that 3DGS has in particular problems with settings where the view rays accumulate matter over a long distance through semi-transparent, yet heterogeneous regions. In such situations, a subtle change of the camera pose can lead to strong changes of the per-pixel accumulated colors and opacities. Thus, 3DGS needs to be optimized for a significantly increased number of parameters, requiring far more Gaussians to accurately represent the data.
@@
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
@@ 4.5 Rendering Performance
[…]
Table 3 shows that the rendering times even on an integrated iGPU is higher than 10 frames per second for the biggest dataset Brain. On current mid- to high-end GPUs, 60 frames per second can be achieved for all datasets. This makes the Cinematic Anatomy pipeline especially appealing for applications where stereoscopic rendering is required. While low memory consumption facilitates efficient rendering on mobile devices, for instance, in mobile AR applications, high rendering performance is required to render two images (one for the left and one for right eye) at sufficient frame rates.
Brain Fullbody Kidney
@@
5 Discussion and Outlook
Our evaluations show that compressed 3DGS enables interactive Cinematic Anatomy with datasets so large that this would have been impossible. This is possible by restricting to static presets. We are confident that this limitation is acceptable for educational use since not more than a few transfer function settings are usually selected. Since the memory requirements of compressed 3DGS are so low, a separate Gaussian representation can be computed for each preset.
In the images we have used so far to perform a qualitative comparison between compressed 3DGS and path tracing, static lighting conditions have been simulated with an environment map that does not change relative to the object. Thus, the object points are seen under the same lighting condition in every view, resulting in rather smooth illumination when changing the perspective under which the viewer looks at the object. This, however, changes when a headlight is used, and a point’s illumination varies with varying camera position. The use of a headlight is demonstrated in Fig. 10. Notably, while most regions can be resolved very well by 3DGS, in some other regions the novel views show reconstruction artifacts. The strong variation of the reflected light under an illumination that changes in every image cannot be captured well by the 3D Gaussian representation. One approach we see to address this limitation is via re-lighting. By generating training images with optical material properties instead of illumination, it might be possible to better recover highly varying lighting conditions at runtime.
An important component in volume rendering applications is an interactively moveable clip plane. Since 3DGS can only handle static scenes, previously unseen objects that become visible due to applying a clip plane (or points that are clipped away and reveal other points) cannot be recovered. Current extensions of 3DGS to 4DGS [57, 57] do not seem applicable in our scenario since they assume that only the positions of visible points change, but not the number of points. Another possibility is to allow only changes of the clip plane in discrete steps, and to compute a separate Gaussian representation for each step. We are confident that a fairly compact representation can be obtained by progressively encoding the object points that appear and disappear when making subsequent steps.
As we have demonstrated, highly transparent volumes cannot be handled very well by 3DGS. Another example demonstrating this is shown in Fig. 11. As long as the volume is homogeneous and doesn’t contain too many interior structures, a quite accurate reconstruction can be achieved. However, we have observed that with increasing depth complexity it becomes more and more difficult for 3DGS to represent all possible color and opacity distributions with a reasonable number of 3D Gaussians. A possible strategy to address this limitation might be to augment 3DGS with 3D Gaussians that are optimized with respect to an object-space loss, similar in spirit to EWA volume splatting [61].
(a) Reconstruction
(b) Ground Truth
Figure 11: Compressed 3DGS with a semi-transparent skull dataset. Fine details are blurred out in the reconstruction.

### v2

@@ 3.3 Compressed Differentiable 3D Gaussian Splatting
Differentiable 3DGS describes an objects by a set of 3D Gaussians
@@
Each Gaussian is centered at $x\in\mathbb{R}^{3}$, and the covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$ describes its orientation and shape. A Gaussian has an opacity $\alpha\in[0,1]$, and a view-dependent color that is represented by a set of spherical harmonics (SH) coefficients.
The 2D projection of a 3D Gaussian is a 2D Gaussian with a covariance that is derived from the view transformation matrix and the Jacobian of the affine approximation of the projective transformation. The scene is rendered by projecting all Gaussian into the image plane in sorted order and blending their contributions.
@@
While Zwicker et al. [ZPVBG01] model a 3D scalar field via a set of 3D Gaussians so that the field can be reconstructed sufficiently well, Kerbl et al. [KKLD23] optimize the position, shape, opacity and SH coefficients of each 3D Gaussian so that their rendering matches a set of initial images of the object. The optimization is performed via differentiable rendering, by taking into account the changes in pixel color due to changes of the 3D Gaussian parameters. The optimization process removes some of the initially selected 3D Gaussians (if no contribution), adaptively splits Gaussians, and modifies their shapes and appearance attributes to minimize an image-based loss function.
To further reduce the memory consumption of 3DGS, we utilize the compression proposed by Niedermayr et al. [NSW24]. It encodes SH coefficients and Gaussian shape parameters into compact codebooks via sensitivity-aware vector quantization, and then fine-tunes the parameters on the training images. Quantization-aware training [RORF16] is used to represent the scene parameters with fewer bits. We call this strategy High-Rate-compression (HR-compression). We also provide an option that uses only quantization-aware training to reduce all scene parameters but the Gaussians’ positions to an 8-bit representation during optimization. We will subsequently call this strategy High-Quality-compression (HQ-compression). Since CA requires the entire data set in focus, we can omit the scaling factor that is usually stored per Gaussian to represent scenes with objects in focus and surrounding background.
@@
Alpha Channel Reconstruction In contrast to classical novel view synthesis, where only RGB colors are reconstructed, in volume rendering applications also the per-pixel accumulated opacity (i.e., alpha) needs to be reconstructed for blending correctly over the background. Therefore, we extend 3DGS to allow for differentiable rendering of images with alpha channel. We use a combination of per pixel L1 and SSIM Loss to accurately reconstruct the alpha channel of the volume rendered training images. As shown in Section 4.4, the reconstruction quality can be improved significantly in this way.
4 Results and evaluation
@@
We analyze the performance, memory consumption and quality of the proposed pipeline for CA with a variety of high-resolution medical data sets showing different anatomical structures. Our 3DGS implementation is a modification of the code provided by Kerbl et al. [KKLD23]. For compression and rendering, we use the settings described by Niedermayr et al. [NSW24].
@@ 4.1 Data Sets
The hierarchical phase-contrast tomography (HiP-CT) data was acquired at the European Synchrotron Radiation Facility (ESRF) in the context of the Human Organ Atlas project[WTW21].
@@
Body is a human CT angiography scan at resolution $317\times 317\times 835$ from the collection by Wasserthal [Was23], image id s0287. The data set contains some semi-transparent material showing significant differences under directional lighting. We render it under complex lighting conditions to challenge 3DGS’s reconstruction capabilities.
Figure 3: From left to right, Brain ($3224\times 3224\times 3585$), Kidney ($1510\times 1706\times 1415$) and Body ($317\times 317\times 835$). All images rendered at full HD with HR-compressed 3D Gaussian splatting using $<$ 70 MB per data set at $>$ 30 frames per second.
We show all data sets in Fig. 3, and provide an interactive online demonstration at https://anonymous-demo-user.github.io/cinematic-3dgs/. For each data set, between one and three presets have been used, including segmentations, transfer function and lighting conditions. 3DGS optimization has been performed on training images of resolution $2048\times 2048$.
@@ 4.2 Preprocessing
With a GPU providing sufficient RAM, the initial images of all data sets can be generated with the publically available CA package, and using the built-in animation system to generate the views. We have used a research version providing batch rendering support, running on an NVIDIA A100 GPU for Brain and an NVIDIA RTX A5000 for Kidney and Body.
Table 1: Memory requirements and preprocessing performance.
Path Tracing 3DGS (HR-Compression)
@@
Body 0.2 GB 23 Min 99 7 MB 50 Min 0.9 M
Table 1 shows in columns Size the size of each data set in GB compared to the size of the final Gaussian representation in MB, when compressed using HR-compression. Column Views shows the number of training images used for differentiable Gaussian splatting optimization. Columns Time show the times to render the initial images via path tracing, and the computation times for generating the compressed Gaussian representations. Note that $90\%$ of the latter time are required by the optimization to generate the 3D Gaussian representation, and only about $10\%$ are consumed by the compression. Column Gaussians gives the number of 3D Gaussians in the final representation. As can be seen in columns Size, the compressed Gaussian representation is so small that it can be downloaded over low-bandwidth channels and rendered on mobile devices equipped with mid- or even low-end GPUs.
@@ 4.3 View Selection
Automatic view selection is demonstrated with Body, which exhibits a lot of structures that are not visible from cameras placed on an ellipsoid around the volume. As a baseline, we reconstruct the volume with images from 256 randomly placed cameras on the ellipsoid. For comparison, we reduce this number to 128 and generate 128 additional cameras with the proposed view selection algorithm (see Fig. 4). As can be seen, overall improved reconstruction quality of parts not seen with random camera selection is achieved.
Figure 4: Automatic view selection generates camera poses covering unseen parts of the volume, like the inside of the rib cage.
@@ 4.4 Quality Evaluation
Fig. 5 compares test images that have not been seen during 3DGS optimization to images rendered with HQ- and HR-compressed 3DGS. Close-up views reveal only subtle color shifts between path traced images and images generated via HR-compressed 3DGS. HQ-compression leads to an increase in memory by a factor of three, yet differences in image quality are further reduced and become so small that they are hardly noticeable by eye.
Notably, significant losses in reconstruction quality are introduced when differentiable 3DGS optimizes only for RGB color (see Fig. 6 for an example). Extending 3DGS so that also opacity is considered in the optimization process improves greatly the reconstruction quality and removes unwanted artifacts caused by the background.
Table 2 shows the average SSIM and PSNR between the test images and the novel views rendered with HR-compressed 3DGS, averaged over all presets. For PSNR and SSIM, only pixels which are not empty (alpha $>0$) in the rendered and ground truth image are considered. PSNR (Alpha) measures the PSNR for the alpha channel between rendered images and ground truths.
Table 2: Quantitative evaluation of HR-compression,
@@
Body 0.87 26.90 29.57
We further shed light on the capabilities of 3DGS to reconstruct semi-transparent regions in a data set. Body is used with a preset so that certain tissue types in the data set become semi-transparent, see Fig. 7. While overall the novel view matches the test images fairly well, the closeup views show that some fine details are not reconstructed accurately, and especially the semi-transparent structures are blurred out. This effect increases with increasing depth complexity, since it becomes more and more difficult for 3DGS to represent all possible color and opacity distributions accurately.
Figure 7: Quality comparison for HQ-compressed and HR-compressed Gaussian representations of semi-transparent regions.
When using a preset with strong directional lighting from the environment map, one observes some high-frequency illumination variations especially in the volumetric regions. This makes it more difficult for 3DGS to accurately recover the tissue structures. Interestingly, Fig. 8 demonstrates that the reconstruction works very well and does not show any severe reconstruction artefacts. At the same time, the semi-transparent regions are again blurred out to a certain extent. We believe that 3DGS has in particular problems with settings where the view rays accumulate matter over a long distance through semi-transparent, yet heterogeneous regions. In such situations, a subtle change of the camera pose can lead to strong changes of the per-pixel accumulated colors and opacities. Thus, 3DGS needs to optimize a significantly increased number of parameters, requiring far more Gaussians to accurately represent the data.
@@ 4.5 Rendering Performance
The WebGPU implementation by Niedermayr et al. [NSW24] is used for performance testing. It enables rendering of compressed 3DGS up to 4×4$\times$4 × faster than the renderer by Kerbl et al. [KKLD23] and runs in a modern browser.
Table 3 shows that the rendering times even on an integrated iGPU is higher than 10 frames per second for the biggest data set Brain. On current mid- to high-end GPUs, 60 frames per second can be achieved for all data sets. This makes the CA pipeline especially appealing for applications where stereoscopic rendering is required. While low memory consumption facilitates efficient rendering on mobile devices, for instance, in mobile AR applications, high rendering performance is required to render two images (one for the left and one for right eye) at sufficient frame rates. In a supplementary video we demonstrate rendering performance of roughly 5 to 20 frames per second on a mobile device with Qualcomm Adreno 740 GPU.
Brain Body Kidney
@@
5 Discussion and Outlook
Our experiments show that compressed 3DGS enables interactive CA with extremely large data sets, by restricting to static presets. We believe that this limitation is acceptable for educational use since not more than a few presets are usually selected. Since the memory requirement of compressed 3DGS is so low, a separate Gaussian representation can be computed for each preset.
In all experiments we have simulated static lighting conditions with an environment map that does not change relative to the object. Thus, the objects are seen under the same lighting condition in every view, resulting in rather smooth illumination when changing the camera pose. This, however, changes when a headlight is used, and a point’s illumination varies with varying camera pose (see Fig. 9). Notably, while in this situation most regions can be resolved very well by 3DGS, in some other regions the novel views show reconstruction artifacts. The strong variation of the reflected light under an illumination that changes in every image cannot be captured well by 3DGS. One approach we see to address this limitation is via re-lighting. By generating training images with optical material properties instead of illumination, it might be possible to better recover highly varying lighting conditions at runtime.
@@ 5 Discussion and Outlook
A useful component for volume exploration is an interactive clip plane, resulting in object points that were previously unseen to become visible. One possibility to include clip planes is to restrict the plane movement to discrete steps and compute a separate Gaussian representation for each step. While this will significantly increase the memory requirements, we are confident that a fairly compact representation can be obtained by exploiting spatial coherence and progressively encoding the 3D Gaussians that appear and disappear when making subsequent steps.